# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v3

@@ IV-E3 Others
There also exist many retrieval augmented works for other video-related tasks. VidIL [348] exploits image-language models to translate video content into temporal-aware prompts with few-shot examples, for various video-language tasks including video captioning, video question answering, video caption retrieval, and video future event prediction. Notably, for trustworthy autonomous driving, RAG-Driver [349] grounds the MLLM in relevant expert demonstrations from a memory database, to produce driving action explanations, justifications, and control signal prediction. Text-to-video generation is to generate video given natural language descriptions.
@@
As shown in Fig. 11, Animate-A-Story [206] develops a framework which can generate high-quality storytelling videos based on texts. It first separates the text into individual plots, and decorates the description using LLM. It then retrieves relevant videos for each plot through a dense retriever [110]. It generates videos through a latent diffusion model, consisting of two branches: a text encoder CLIP, and a structure encoder which takes the estimated depth of the retrieved videos as structure control.
@@ IV-F1 Audio Generation
The goal of audio generation is to generate audio with natural language input.
@@
Re-AudioLDM [159] adopts dense retriever CLAP [26] to retrieve similar caption-audio pairs given input prompt. As shown in Fig. 12, the generator, latent diffusion model and VAE-based decoder, take the representations of input text and retrieved pairs as input and generate output audio. Make-An-Audio [44] uses dense retriever CLAP [26] to augment data, retrieving related audio given natural language text. It then constructs pseudo prompts for diffusion-based text-to-audio model training.
@@ IV-F2 Audio Captioning
The goal of audio captioning is to generate natural language data with audio data, which is basically a sequence-to-sequence task. RECAP [350] leverages CLAP [26] to retrieve related captions given audio data. The retrieved captions are then included into the prompt input for GPT-2 model, which interacts with audio embeddings through cross attention. In [43], dense retriver VGGish [107] is adopted to produce dense embedding of audio data, and GPT-2 is adopted to generate representations of the retrieved captions which are paired with similar audios. After obtaining the representations of audios and captions, an extra multi-head attention block and a linear layer fuses all the information and generates the output. Some research studies transform audio modality to text, in order to leverage advancements in LLMs [351, 352, 353]. They take advantage of deep retrieval models, aligning the modalities into the same latent space for downstream text generation.
@@ IV-G1 Text-to-3D
Retrieval can be applied to augment the generation of 3D contents. ReMoDiffuse [51] aims at generating motions using diffusion models. It first retrieves relevant motion entites through CLIP given text input, then leverages the information of the text and the retrieved entities through a semantic-modulated attention layer.
@@
AMD [158] designs two branches of motion diffusion for fidelity and diversity. As shown in Fig. 13, the first branch inputs the original prompt text for diffusion; the second branch decomposes the input text into anatomical scripts and retrieve similar reference motions for diffusion. A transformer-based fusion module is further applied to adaptively balance the result from two branches. RetDream [50] targets general 3D generation, using retrieved 3D assets to augment the process of variational score distillation from 2D diffusion models. Given an input query, it retrieves relevant 3D assets through CLIP, then utilizes the retrieved assets to provide geometric prior and adapted 2D prior. Concretely, retrieved assets not only impose an additional velocity on particles for distribution initialization, but also help optimize the 2D diffusion model through low-rank adaptation.
@@ V Benchmark
Chen et al. [361] proposed an RAG benchmark, which evaluates RAG from four aspects: Noise Robustness, Negative Rejection, Information Integration, and Counterfactual Robustness, respectively. Noise Robustness evaluates whether LLMs could extract the necessary information from documents containing noisy information. The noisy information is relevant to the input query but useless for answering it. Negative Rejection measures whether LLMs would reject to respond the query when the retrieved content is not enough. Information Integration assesses whether LLMs could acquire knowledge and make responses by integrating multiple retrieved contents. Counterfactual Robustness refers to the ability of LLMs to identify counterfactual errors in the retrieved content.
Another three benchmarks, RAGAS [362], ARES [363] and TruLens [364], consider three different aspects: Faithfulness, Answer Relevance, and Context Relevance, respectively. Faithfulness focuses on the factual errors in the results when the correct answers can be inferred from the retrieved contents. Answer Relevance measures whether the generated results actually address the problems (i.e., queries) or not. Context Relevance judges whether the retrieved contents contain as much knowledge as possible to answer the queries, and as little irrelevant information as possible.
CRUD-RAG [365] divides all RAG tasks into four categories, which are Create, Read, Update, and Delete, respectively, and also evaluates each category using text continuation, question answering (with single- and multi-document questions), hallucination modification, and open-domain multi-document summary. MIRAGE [366] is a benchmark designed for assessing the application of RAG in the medical domain, focusing on the comparison and optimization of medical question-answering systems’ performance. KILT [367] is another benchmark focuses on ensuring information accuracy and reliability by aligning Wikipedia pages with specific snapshots and pinpointing the most pertinent text ranges through BLEU score evaluations. It filters out lower-quality data to maintain a high standard of information mapping, offering a variety of retrieval system options like TF-IDF, DPR, RAG, and BLINK + flair to support evidence-based predictions or citations according to task requirements.
@@ VI-A Limitations
Despite the widespread adoption of RAG, it suffers from several limitations by nature. In this paper, we provide a summary of the limitations and engage in an in-depth discussion.
@@ VI-A1 Noises in Retrieval Results
Information retrieval cannot yield perfect results because information loss appears in representations generated by encoder models. Additionally, ANN search can also provide approximate results rather than exact ones. Consequently, certain degree of noise is inevitable in retrieval results, manifesting as irrelevant objects or misleading information, which may cause failure points in RAG systems [368]. Though the common sense is that increasing the accuracy of retrieval will contribute to the effectiveness of RAG, a recent study surprisingly shows that noisy retrieval results may conversely help improve the generation quality [369]. A possible explanation is that diversity in retrieval results may also be necessary for prompt construction [370]. As a result, the impact of noise in retrieval results remains uncertain, leading to confusion in practical uses regarding which metric to employ for retrieval and how to facilitate the interaction between the retriever and the generator.
@@ VI-A2 Extra Overhead
While retrieval can help mitigate the costs of generation in certain cases [30, 31, 32], the incorporation of retrieval sometimes introduces non-negligible overhead. Considering that RAG is primarily employed to improve the performance of existing generative models, the inclusion of additional retrieval and interaction processes leads to increased latency. Worse still, when combined with complex enhancement methods, such as recursive retrieval [371] and iterative RAG [218], the extra overhead will become even more significant. Furthermore, as the scale of retrieval expands, the storage and access complexity associated with data sources will also increase. In presence, RAG systems exhibit a trade-off between costs and benefits. Looking ahead, we anticipate further optimization to alleviate the associated overhead [372].
@@ VI-A3 The Gap between Retrievers and Generators
Seamlessly integrating retrieval and generation components requires meticulous design and optimization. Since the objectives of retrievers and generators may not align, and their latent spaces might differ, designing their interaction poses challenges. As introduced in Section III, numerous approaches have been proposed to enable effective RAG, and these approaches either disentangle the retrieval and generation processes or integrate them at an intermediate stage. While the former is more modularized, the latter could potentially benefit from joint training. Till not, there lacks a sufficient comparison of different ways of interaction across various scenarios.
@@ VI-A4 Increased System Complexity
The introduction of retrieval unavoidably increases the system complexity and the number of hyper-parameters to tune. For instance, a recent study on the trade-off between attribution and fluency in prompt-augmentation-style RAG demonstrates that using top-k retrieval for generation improves attribution, but hurts fluency in turns [373]. The counter effects of different aspects in RAG, such as metric selection, are still under explored. Therefore, further refinement of RAG systems, both in terms of algorithms, and deployment, is necessary to fully unlock their potentials.
@@ VI-B Potential Future Directions
VI-B1 More Advanced Research on RAG Methodologies, Enhancements, and Applications
A straight-forward research direction is to develop more advanced methodologies, enhancements, and applications of RAG.
@@ VI-B1 More Advanced Research on RAG Methodologies, Enhancements, and Applications
As introduced in Section III-A, existing works have explored various interaction patterns between retrievers and generators. However, since the optimization target of these two components are distinct, the practical augmentation process has a large impact on the final generation results. Investigation of more advanced foundations for augmentation holds promise for fully unleashing the potential of RAG.
Based on a constructed RAG system, enhancements are helpful to improve the effectiveness of certain components or the entire pipeline. Given the inherent complexity of the system, there exists significant potential for RAG to improve, necessitating proper tuning and careful engineering. We look forward to further experimental analysis and in-depth exploration that will contribute to the development of more effective and more robust RAG systems.
As introduced in Section IV, RAG is a general technique that has been applied across diverse modalities and tasks. Yet most of existing works straightforwardly integrate external knowledge with the specific generation tasks, without thoroughly taking into account the key characteristics of the target domains. Therefore, for generation tasks that do not fully leverage the power of RAG, we are confident that designing proper RAG system will be beneficial.
VI-B2 Efficient Deployment and Processing
@@
Currently, several deployment solutions of query-based RAG for LLMs have been proposed, such as LangChain [375], LLAMA-Index [175], and PipeRAG [376]. However, for other foundations of RAG and/or generation tasks, there lacks a plug-and-play solution. In addition, given the extra overhead introduced by retrieval, and considering that the complexities of both the retriever and generator will continue to grow, achieving efficient processing in RAG remains a challenge, necessitating targeted system optimization.
VI-B3 Incorporating Long-tail and Real-time Knowledge
@@
While a key motivation of RAG is to harness real-time and long-tail knowledge, few studies have explored the pipeline for knowledge updating and expansion. Many existing works make up the retrieval sources with merely the training data of generators, thereby neglecting the dynamic and flexible information advantages that could have been offered by retrieval. As a consequence, designing a useful RAG system with continuously updated knowledge and/or flexible knowledge sources, along with corresponding system-level optimizations, is a growing research direction. With the capability of utilizing long-tail knowledge, we also expect RAG to leverage personalized information and features, so as to adapt to today’s web service.
VI-B4 Combined with Other Techniques
@@
In essential, RAG is orthogonal to other techniques that share the goal of improving AIGC effectiveness, including fine tuning, reinforcement learning, chain-of-thought, agent-based generation, and other potential optimizations. However, the exploration of simultaneously applying these techniques is still in its early stages, calling for further research to delve into algorithm design and fully leverage their potential. It is worthy to note that a recent notion appears “long-context models like Gemini 1.5 will replace RAG”. Nevertheless, this assertion does not hold true — RAG exhibits greater flexibility in managing dynamic information, encompassing both up-to-date and long-tail knowledge [377]. We believe that RAG in the future will take advantage of long context generation to achieve even better performance, rather than simply being weeded out by it.

### v4

@@ IV-E3 Others
RAG also works for other video-related tasks. VidIL [304] converts video content into temporal-aware LLM prompts for tasks like video captioning, question answering, and future event prediction.
@@
For trustworthy autonomous driving, RAG-Driver [305] grounds the MLLM in retrieved expert demonstrations, to produce driving action explanations. Animate-A-Story [163] simplifies text-to-video generation by dividing it into plot-based video augmentation and video-diffusion generation conditioned on text and video inputs.
@@ IV-F1 Audio Generation
Audio generation usually synthesises audio with natural language prompt.
@@
Given input prompt, Re-AudioLDM [115] retrieves relevant caption-audio pairs with dense retriever CLAP [26] for generation. Make-An-Audio [44] retrieves audios given text prompt, then constructs pseudo prompts for text-to-audio diffusion model training.
@@ IV-F2 Audio Captioning
Audio captioning, basically a sequence-to-sequence task, generates natural language data for audio data. RECAP [306] and [43] leverages dense retrievers, CLAP [26] and VGGish [68] respectively, to retrieve related captions given audio data. For RECAP, captions are included into LLM prompts, while  [43] uses both audio and retrieved captions in attention module. Other research studies align audio modality with text to leverage advancements in LLMs [307, 308, 309] for various downstream text generation.
@@ IV-G1 Text-to-3D
Retrieval can be applied to augment 3D asset generation.
@@
ReMoDiffuse [51] retrieves relevant motion entities and generates motions using diffusion models, with the semantic-modulated attention and condition mixture guidance. AMD [114] designs and fuses two motion diffusion models. One branch conditions on the original prompt, while the other decomposes the prompt into anatomical scripts and retrieves similar motions. RetDream [50] retrieves 3D assets to augment the variational score distillation [310] of 2D diffusion models. These assets offer geometric and adapted 2D priors, which not only impose additional velocity on particles for initialization but also help optimize 2D diffusion models by LoRA.
@@ V Benchmark
Chen et al. [318] proposed an RAG benchmark that evaluates across four dimensions: (1) Noise Robustness, testing if LLMs can extract necessary information from noisy documents; (2) Negative Rejection, assessing if LLMs can reject to respond when retrieved content is insufficient; (3) Information Integration,checking if LLMs can acquire knowledge and respond by integrating multiple retrieved contents; (4) Counterfactual Robustness, determining if LLMs can identify counterfactual errors in retrieved content.
Another three benchmark, RAGAS [319], ARES [320] and TruLens [321], consider another three aspects with the help of a separate evaluator LLM: (1) Faithfulness, which examines factual errors in results when correct answers can be inferred from the retrieved contents; (2) Answer Relevance, which measures whether generated results actually address the queries; (3) Context Relevance, which judges whether retrieved content provides relevant knowledge while minimizing irrelevant information.
CRUD-RAG [322] categorizes RAG tasks into four types: Create, Read, Update, and Delete. It evaluates each category using text continuation, question answering (including single- and multi-document questions), hallucination modification, and open-domain multi-document summarization. MIRAGE [323] assesses RAG in the medical domain, focusing on the performance of medical question-answering systems. KILT [324] aligns Wikipedia snapshots to verify information accuracy, using BLEU scores to pinpoint relevant texts and filtering to uphold quality, thus providing diverse retrieval systems for evidence-backed predictions or citations.
@@ VI-A Limitations
Despite the widespread adoption of RAG, it suffers from several limitations by nature.
@@ VI-A1 Noises in Retrieval Results
Information retrieval is inherently flawed due to information loss in item representations and ANN search. The inevitable noise, manifesting as irrelevant content or misleading information, can create failure points in RAG systems [325]. However, although improving retrieval accuracy seems intuitive for RAG effectiveness, recent research surprisingly finds that noisy retrieval results might enhance generation quality [326]. A possible explanation is that diverse retrieval outcomes could contribute to prompt construction [327]. Thus, the impact of retrieval noise remains unclear, leading to confusion about metric selection and retriever-generator interaction in practical uses.
@@ VI-A2 Extra Overhead
While retrieval can reduce generation costs in certain cases [30, 31, 32], it incurs non-negligible overhead in most cases. In other words, the retrieval and interaction processes increase latency inevitably. This is amplified when RAG is combined with complex enhancement methods, such as recursive retrieval [328] and iterative RAG [175]. Furthermore, as the scale of retrieval sources expands, the storage and access complexity will also increase [329]. Such overhead hampers the practicality of RAG in real-time services that are sensitive to latency.
@@ VI-A3 The Gap between Retrievers and Generators
Since the objectives of retrievers and generators may not align, and their latent spaces might differ, designing their interaction requires meticulous design and optimization. Current approaches either disentangle retrieval and generation or integrate them at an intermediate stage. While the former is more modular, the latter could benefit from joint training but hamper generality. Selecting a cost-effective interaction method to bridge the gap poses a challenge and necessities deliberation in practice.
@@ VI-A4 Increased System Complexity
The introduction of retrieval unavoidably increases the system complexity and the number of hyper-parameters to tune. For instance, a recent study found that using top-$k$ rather than a single retrieval improves attribution but harms fluency in query-based RAG [330], while other aspects such as metric selection are still under explored. Thus, it requires more expertise to tune the generation service when RAG is involved.
@@ VI-B Potential Future Directions
VI-B1 Novel Design of Augmentation Methodologies
@@
Existing research has explored various interaction patterns between retrievers and generators. However, due to distinct objectives in these two components, the practical augmentation process has a significant impact on the final generation results. Investigation of more advanced foundations for augmentation holds promise for fully unleashing the potential of RAG.
VI-B2 Flexible RAG Pipelines
RAG systems are progressively embracing flexible pipelines, such as recursive, adaptive, and iterative RAG. Alongside proper tuning and careful engineering of each component, the unique combination of retrieval sources, retrievers, generators, and RAG subsystems holds promise for handling complex tasks and enhancing overall performance. We eagerly anticipate pioneering exploration that will drive the evolution of even more innovative RAG systems.
VI-B3 Broader Applications
RAG is a general technique applied in various applications. However, some generative tasks have not yet explored RAG, and in many domains, RAG is applied naively without considering the domain’s unique characteristics. We believe designing domain-specific RAG techniques will significantly benefit broader applications.
VI-B4 Efficient Deployment and Processing
@@
There exist several deployment solutions for query-based RAG with LLMs, such as LangChain [332], LLAMA-Index [131], and PipeRAG [333]. However, for other RAG foundations and/or generation tasks, there lacks a plug-and-play solution. Besides, due to retrieval overhead and increasing complexities in retrievers and generators, achieving efficient RAG is still challenging and necessitates further system-level optimizations.
VI-B5 Incorporating Long-tail and Real-time Knowledge
@@
While a key motivation of RAG is to harness real-time and long-tail knowledge, few studies have explored the pipeline for knowledge updating and expansion. Many existing works use merely the generators’ training data as retrieval sources, neglecting the dynamic and flexible information that retrieval could offer. As a consequence, there is a growing research on designing RAG systems with continuously updated knowledge and flexible sources. We also expect RAG to step further, adapting to personalized information in today’s web service.
VI-B6 Combined with Other Techniques
@@
RAG is orthogonal to other techniques that also aim to improve AIGC effectiveness, such as fine-tuning, reinforcement learning, chain-of-thought, and agent-based generation. The combining of these methods is still in its early stages, calling for further research to fully exploit their potential through novel algorithm designs. It is worthy to note that a recent notion appears “long-context models like Gemini 1.5 will replace RAG”. Nevertheless, this assertion overlooks RAG’s flexibility in managing dynamic information, encompassing both up-to-date and long-tail knowledge [334]. We expect RAG to benefit from long context generation, rather than being replaced by it.